OpenStudy (anonymous):

Graph the linear inequality: y<1/9x+9

ganeshie8 (ganeshie8):

start by graphing the corresponding "line" : y = 1/9x + 9

ganeshie8 (ganeshie8):

do u knw how to graph a line ?

OpenStudy (anonymous):

yes, but how do I graph that?

OpenStudy (anonymous):

is this the y=mx+b thing? lol

ganeshie8 (ganeshie8):

exactly ! first graph the line...

ganeshie8 (ganeshie8):

y = 1/9x + 9 it cuts the y-axis at "9"

OpenStudy (anonymous):

b is y, right? so it'll be (0,9) then the slope is 1/9?

ganeshie8 (ganeshie8):